OpenStudy (anonymous):

Given the linear equation y=1/4 x-2, find the y-coordinates of the points (-8, ), (-4, ), and (4, ). Please show all of your work. Plot those points and graph the linear equation. Submit your graph through the dropbox.

OpenStudy (anonymous):

In all given coordinates, for example, if you were given a coordinate (4,5) the first value (4) is the x-value, and the second value (5) is the y-value. So in the case of (-8, ) you are given the x-value, but you need the y-value. So you would go back to the original equation y=1/4 x - 2 and plug in -8 for x so you get: y= 1/4(-8) - 2 then y= -7.75 - 2 so y= -9.75 in the first problem. Then you apply the same rules to the other two problems.

OpenStudy (anonymous):

I am very lost sorry